OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 2 and a y-intercept of -3.

OpenStudy (anonymous):

@Shadowys @ash2326 @Agent_Sniffles hellppp

OpenStudy (anonymous):

@ganeshie8 can you help any one plzz

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

slope intercept form of a straight line is \[y=mx+b\] where \(m\) is the gradient and \(b\) is the \(y\)-intercept

OpenStudy (unklerhaukus):

\[\color{gray}{\textrm{gradient means slope}}\]

OpenStudy (anonymous):

but i dont understand how to do the problem

OpenStudy (unklerhaukus):

read the question

OpenStudy (anonymous):

so its 2m - 3 x

OpenStudy (unklerhaukus):

an equation is a mathematical statement with an equals sign

OpenStudy (anonymous):

never mind im lost never mind

OpenStudy (anonymous):

y=-3+2 correct or not

OpenStudy (unklerhaukus):

you need an \(x\) in the some place

OpenStudy (unklerhaukus):

there*

OpenStudy (anonymous):

yx=-3+2

OpenStudy (unklerhaukus):

the slope intercept form of a straight line is \[y=mx+b\] find \(m\) & find \(b\) from the question

OpenStudy (anonymous):

i dont know what m and be are there are olny 2 number 2 and -.3 so im guessing its one of thoses

OpenStudy (unklerhaukus):

a slope of 2 \(\Longrightarrow\) m=2 a y-intercept of -3. \(\Longrightarrow\) b=-3

OpenStudy (anonymous):

so its y=2x+-3

OpenStudy (anonymous):

correct or not

OpenStudy (anonymous):

r u still there

OpenStudy (unklerhaukus):

yeah thats rights

OpenStudy (anonymous):

thank you

OpenStudy (unklerhaukus):

you can simplify +- to -

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3x-2y=6